Find two consecutive even integers if the square of the smaller integer is 10 more than the larger integer.
Answer by mananth(16949)

let the integers be x & x+2
x^2= 10+x+2
x^2=x+12
x^2-x-12=0
(x-4)(x+3)=0
x=4,x=-3
4, 6